OpenStudy (anonymous):

35 is what percent of 70

OpenStudy (anonymous):

\[\frac{ 35 }{ 70 } \times 100 percent=50 per cent\]

OpenStudy (texaschic101):

35 is what % of 70 " of " means multiply and " is " means equals. 35 = (x%)(70) divide by 70 to get x% by itself 35/70 = x% .5 = x% Now turn .5 into a percent by moving the decimal two spaces to the right. 50% = x check... 35 = (.50)(70) 35 = 35 (correct) ANSWER : 50%